# Machine Learning with Physics Knowledge for Prediction: A Survey

### v1

@@ 3.1 Learning Differential Models from Data
A challenge in learning models of physical processes is that our reality exists in continuous time, while the numerical algorithms require discretization. A useful inductive bias in this setting is to define a (black-box) model in continuous time by incorporating integration into prediction. The result is a more flexible predictive model that can learn and forecast on irregular time intervals rather than a rigid discretization. This approach is referred to as learning the differential equations.
@@
The connection between ODEs and deep learning, neural ODE, is a powerful tool. Perhaps the earliest endeavor of learning differential equations is the ResNet architecture (i.e., residual flow block) proposed by He et al. (2016), although the original motivation was to mitigate gradient attenuation for parameters in early layers. A ResNet block defines a discrete-time residual transformation
@@
The common optimization paradigm in learning neural function approximations is backpropagation through learning objectives. When learning ODEs, the objective is generally to fit the parameters ${\bm{\theta}}$ of ${\bm{f}}$ such that the flow gradient at ${\bm{x}}$ matches the reference transform at time $t$ (Kidger, 2022). This paradigm typically requires a sufficiently large dataset to avoid overfitting for learning highly non-linear ODEs; otherwise, one should resort to classical (Voss et al., 2004) or neural (Chen et al., 2021; Forgione and Piga, 2021) system identification methods. To obtain the gradient of the objective w.r.t. the parameters ${\bm{\theta}}$, it is required to propagate the gradients through the numerical integration. A straightforward approach is to store all of the function evaluations in the forward pass of the ODE and backpropagate through the whole numerical integration. Yet, the space complexity increases linearly with time $T$, which becomes a computational burden for long-time series and high-order numerical solvers (Norcliffe and Deisenroth, 2023). Given a scalar-valued loss function $L$, the adjoint of ${\bm{x}}(t)$, ${\bm{a}}(t)={\textnormal{d}}L/{\textnormal{d}}{\bm{x}}(t)$, having the dynamic derived from the chain rule,
@@
As the gradient is obtained as the solution of the time-reversed ODE, the gradients can be evaluated without storing any additional data. Thus, the adjoint method optimizes computational resources by balancing time against space complexity. Several works (Gholaminejad et al., 2019; Zhuang et al., 2020; Onken and Ruthotto, 2020) showed that the gradient estimation based on the adjoint method is error-prone due to the numerical inaccuracies of solving the backward ODE, due to the typical high Lyapunov characteristic exponent of highly non-linear ODEs. Many improved backward integrators have been proposed to mitigate this issue (Zhuang et al., 2020; 2021a; Gholaminejad et al., 2019).
@@
This can prevent the trajectories of the system from intersecting each other and thus allows the representation of more general system behaviors. Furthermore, according to Dupont et al. (2019), augmented neural ODEs can achieve lower losses, better generalization, and lower computational cost than regular neural ODEs, also since time discretizations can be coarser as the learned rate functions ${\bm{f}}_{\bm{\theta}}$ are smoother. However, challenges are the determination of the size of ${\bm{h}}$ and training without knowing the values of ${\bm{h}}$.
@@
Earliest works propose to learn a convolutional neural network as a map between finite Euclidean spaces $f_{{\bm{\theta}}}$ (Guo et al., 2016; Zhu and Zabaras, 2018; Bhatnagar et al., 2019). Building on these approaches, auto-regressive methods have been introduced to estimate the solution at discrete time steps ${{\bm{u}}({\bm{x}},t+\Delta t)=f_{\bm{\theta}}(\Delta t,{\bm{u}}({\bm{x}},t))}$, and either, mix with conventional numerical methods in classical PDE solvers (Bar-Sinai et al., 2019; Greenfeld et al., 2019; Hsieh et al., 2019), or propagate the solution through time by recurrent calls of the networks (Brandstetter et al., 2022). All previous learning PDE methods have in common that they rely on a discretization of the domain, yet in many scenarios, it is desirable to evaluate the solution at any point in the domain. For these reasons, learning approaches that directly learn a mapping from a point in the domain to its solution $u_{{\bm{\theta}}}$ have been actively explored in recent years (Yu et al., 2018; Raissi et al., 2019; Bar and Sochen, 2019), which introduces physics constraints as losses, is covered in detail in section 3.2. Yet, these models come at the cost that they provide solutions for a particular PDE instance and do not generalize to different inputs. Addressing these contemporary issues of learning PDEs, directly learning the solution operator of the PDE as a map between infinite-dimensional spaces has gained traction in recent years (Lu et al., 2019; Li et al., 2020a; 2021a). These models extend the fundamental work of Chen and Chen (1995) that first showed a universal approximation theorem for learning operator networks. These models are explored in more detail in section 3.4.
3.2 Learning Solution Fields with Physics-Informed Losses
where ${\bm{v}}_{m}$ are test functions and $M$ is the number of sampling points. This approach is attractive because it can lower the order of differential operators that the neural network approximates, making the computation of loss functions cheaper. However, the analytic tractability puts severe restrictions on the models and settings that can be considered, and the authors only demonstrate the approach on a one- and two-dimensional case of Poission’s equation.
The practical challenges of applying PINNs in dynamical systems lie in approximating and optimizing the loss function. Due to the nature of differential equations, the residual loss is often not ’fixed’ between time steps, as it requires the model to be self-consistent with a set of its own derivatives. This objective means the regression target is not fixed, and the model is regressed against itself during optimization until self-consistency is achieved. This ‘bootstrapping’ of a function approximator can lead to unstable optimization in practice [^2] Bootstrapping and function approximation comprises of two parts of the ‘deadly triad’ in reinforcement learning (Sutton and Barto, 2018) where this model bootstrapping is also performed when computing value functions. As a result, sampling and optimization must be designed carefully and jointly to ensure successful optimization.
Optimization. Wang et al. (2021a) investigate the training dynamics of PINNs and observed that the magnitude of the parameter gradient increased during training, in part due to the multi-objective loss function. To achieve stable convergence, the authors advocated for a learning rate schedule.
Fonseca et al. (2023) studied the optimization landscape of PINNs, and found that their solutions lie in sharp minima, i.e., regions of the loss landscape where the loss varies significantly under small perturbations of the solution. This is in contrast to folklore in deep learning, where solutions in flat minima are preferred as they are interpreted as solutions that are not overfitting and, therefore, generalizing. Since PINNs models are trained in their domain, the issue of out-of-distribution generalization is less of a concern. This difference results in typical deep learning optimization techniques being less desirable. The authors found that quasi-Newton solvers such as the limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm (LBFGS) were more effective at minimizing the PINNs objective.
@@ 3.3 Learning Models with Algebraic Structures
where ${\bm{\tau}}$ is the generalized force. Liu et al. (2005) optimize directly the system trajectory and dynamic parameters, minimizing the violation of the constraint from equation  (17). Much later on, by realizing the rigid-body kinetic energy ${\mathcal{T}}({\bm{q}},\dot{{\bm{q}}})=\frac{1}{2}\dot{{\bm{q}}}^{\intercal}{\bm{M}}({\bm{q}})\,\dot{{\bm{q}}}$ under gravitational potential field ${\textnormal{d}}{\mathcal{V}}/{\textnormal{d}}{\bm{q}}={\bm{g}}({\bm{q}})$, Lutter and Peters (2023) propose deep Lagrangian networks (DeLAN), parameterizing directly the inertia tensor ${\bm{M}}_{{\bm{\theta}}}({\bm{q}})$ with positivity constraint ensuring dynamical stability property, and the gravity force ${\bm{g}}_{{\bm{\theta}}}({\bm{q}})$ (Figure 3). Combining automatic differentiation with the state partial derivatives, the derived derivatives w.r.t. time and, thus, the violation of Lagrangian constraint are efficiently computed in a forward pass (see Figure 3), effectively facilitating robot model learning with sample efficiency. Cranmer et al. (2020b) then generalizes DeLAN to the Lagrangian Neural Network (LNN), which assumes a more general form of kinetic energy rather than rigid-body kinetics. The LNN works by directly computing the matrix inverse of the Hessian of a learned Lagrangian $\left(\frac{\partial^{2}{\mathcal{L}}}{\partial\dot{{\bm{q}}}^{\top}\partial\dot{{\bm{q}}}}\right)$, and computing the implied acceleration with:
@@
where the inverse is typically approximated with a Moore-Penrose pseudoinverse. Training is then done via backpropagating through a loss between the true and predicted $\ddot{{\bm{q}}}$.
@@
with neural networks on the damping ${\bm{D}}$ matrix, the Hamiltonian ${\mathcal{H}}$, the generalized force ${\bm{G}}({\bm{q}}){\bm{u}}$, effectively recovering the underlying stationary Hamiltonian, time-dependent force, and dissipative coefficient.
@@
where ${\bm{q}}(t)\in{\mathbb{R}}^{d}$ are the dynamic states and ${\bm{z}}(t)\in{\mathbb{R}}^{n}$ are the algebraic variables. This class of differential-algebraic equations has been used to model various engineering applications, such as power network systems with frequency-dependent dynamic loads or nonlinear oscillations. Under the AINN framework, Moya and Lin (2023) model $f,g$ as neural networks and treat both statements in equation  (22) as algebraic relations. Given a set of training data on ${\bm{q}}$, they perform a model rollout with an implicit Runge-Kutta method (Iserles, 2009) from the initial conditions, then learn $f,g$ by matching the rollout trajectory with data and minimizing the violation of the constraint $g({\bm{q}},{\bm{z}})$, satisfying the defined algebraic relations.
AINN is still in its infancy state with very few works. However, we foresee vast applications of parametrizing algebraic rules with neural networks in many hybrid systems (i.e., discrete-continuous) in robotics or engineering applications.
@@ 3.4 Operator Learning
So far in this survey, we have considered learning a single solution $u_{\bm{\theta}}({\bm{x}})$ to a differential equation, which requires solving the dynamical system over the domain of interest for a given set of system parameters, such as boundary conditions. A more general model to learn would be the mapping from system parameters to solution, which could enable extrapolating to new system parameters without having to solve the system. This is the idea behind learning operators. When the system parameters are expressed as a state-varying function ${\bm{a}}({\bm{x}}):\Omega\rightarrow{\mathbb{R}}^{d_{a}}$, this operator maps from the infinite-dimensional function-space input to the corresponding solution ${\mathsf{L}}_{\bm{a}}^{-1}{\bm{f}}({\bm{x}})={\bm{u}}_{\bm{a}}({\bm{x}})$, similarly a function-space output (Figure 4). Operating on functions introduces the same flexible discretization invariance discussed in section 3.1, but now extended beyond time to any continuous state space. Moreover, the training data for such models now extends across several instances of the physical system, so the model is extended in a physics-informed fashion to be conditioned on the system parameters. We will revisit multi-instance learning in section 4 in the context of learning data-driven priors.
@@
Kovachki et al. (2023) report that training with the relative $L^{2}$ loss equation  (23) is less prone to overfitting due to normalization compared to the alternative mean squared error (MSE) loss. The specific parameterization of the operator networks has been the primary focus of recent approaches to guarantee discretization invariance, which has a bounded generalization error while being computationally efficient. Following, we will discuss the two prevailing operator learning architectures, DeepONets (Lu et al., 2019) and neural operators (Li et al., 2020a) highlighted in Figure 4. Although both architectures follow universal approximation theorems to learn arbitrary operators, their conceptual design is fundamentally different.
Figure 4: (Left) A schematic of operator networks that learn a functional mapping from the inputs $a$ to the PDE solution ${\bm{u}}$ at a designated point ${\bm{x}}$. (Middle) DeepONets can approximate any non-linear operator by learning a finite-dimensional mapping $\hat{\bm{u}}={\bm{\varphi}}_{\bm{\theta}}(\hat{\bm{a}})$ in a latent space which is spanned by linear encodings $\hat{\bm{a}}={\bm{e}}_{\bm{\theta}}({\bm{a}})$ and linear decoding ${\bm{u}}={\bm{g}}_{\bm{\theta}}(\hat{\bm{u}},{\bm{x}})$. (Right) Neural operators directly map a discretized representation of the input ${\bm{a}}_{1:N}$ to the solution at an arbitrary discretization ${\bm{u}}_{1:M}$.
@@
In order to deal with infinite-dimensional input and output spaces, deep operator networks (DeepONets) (Lu et al., 2019) are based on the idea of learning the solution operator in a finite-dimensional space ${\bm{\varphi}}:{\mathbb{R}}^{d_{{\bm{e}}}}\rightarrow{\mathbb{R}}^{d_{{\bm{g}}}}$. This requires a functional mapping of the input functions ${\bm{a}}$ into a latent representation ${\bm{e}}:{\mathcal{A}}\rightarrow{\mathbb{R}}^{d_{{\bm{e}}}}$ and a respective functional mapping ${\bm{g}}:{\mathbb{R}}^{d_{{\bm{g}}}}\rightarrow{\mathcal{U}}$ that decodes. Thus, the operator learning architecture based on the finite-dimensional mappings can be formulated as ${\mathsf{S}}_{\bm{\theta}}{\bm{a}}={\bm{g}}_{\bm{\theta}}\circ{\bm{\varphi}}_{\bm{\theta}}\circ{\bm{e}}_{\bm{\theta}}$. For this architecture, Chen and Chen (1995) first showed that for linear maps ${\bm{e}}_{\bm{\theta}}$ and ${\bm{g}}_{\bm{\theta}}$, and a single-layer network ${\bm{\varphi}}_{\bm{\theta}}$, there exist finite-dimensional latent spaces ${\mathbb{R}}^{d_{\bm{e}}}$ and ${\mathbb{R}}^{d_{\bm{g}}}$ for which any non-linear operator can be approximated. This universal approximation theorem for operator learning has recently been expanded to encompass arbitrary deep-learning architectures for mapping in finite-dimensional spaces (Lu et al., 2019). The different embodiments of the three modules lead to distinct approaches, all underlying the same concept of finite-dimensional mappings.
@@
Similar ideas within the framework of reduced-order methods have been presented by Peherstorfer and Willcox (2016). This method, known as operator inference, involves projecting the state-space onto a lower-dimensional representation using proper orthogonal decomposition (POD) and subsequently learning the linear operators of systems of nonlinear ordinary differential equations (ODEs). However, these models often lack the generalization capabilities seen in more recent methodologies such as neural operators.
@@
In numerous relevant applications, the PDE operator exhibits non-linearity, rendering equation  (24) inapplicable. Although approaches utilizing encoding and decoding schemes to transform the problem into sub-spaces amenable to linear operators have been suggested (Gin et al., 2021), they lack the theoretical assurances inherent in operator learning methods. Yet, the idea of using integral kernel operators play a pivotal role in designing neural operators with theoretical guarantees
@@
The kernel integration layers are composed of a linear operator, represented by the matrix ${\bm{W}}_{t}\in{\mathbb{R}}^{d_{v_{t}}\times d_{v_{t-1}}}$, a bias term ${\bm{b}}_{t}\in{\mathbb{R}}^{d_{v_{t}}}$, and the integral kernel operator (equation  (25)). By parameterizing ${\mathsf{P}},{\mathsf{Q}},{\bm{W}}_{t},{\bm{b}}_{t},\kappa_{t}$, the neural operator can be shown to approximate arbitrary non-linear operators (Kovachki et al., 2023). Although NOs are not restricted to finite-dimensional mappings, as noted by (Kovachki et al., 2024), in practice, NOs work on discretized representations of the input domain ${\bm{a}}_{1:N}$. The main reason is that the integral kernel operator (equation  (25)) remains a computationally demanding operation. Therefore, we will examine different architectures to approximate ${\mathsf{K}}$ efficiently.

### v2

@@ 3.1 Learning Differential Models from Data
A challenge in learning models of physical processes is that our reality exists in continuous time, while the numerical algorithms require discretization. A useful inductive bias in this setting is to define a black-box model in continuous time by incorporating integration into prediction. The result is a more flexible predictive model that can learn and forecast on irregular time intervals rather than a rigid discretization. This approach is referred to as learning the differential equations.
@@
The connection between ODEs and deep learning, the neural ODE, enjoys a long history. Perhaps the earliest predecessor to ‘deep’ learning of differential equations from data is the ResNet architecture (i.e., residual flow block) proposed by He et al. (2016), although the original motivation was to mitigate gradient attenuation for parameters in early layers. A ResNet block defines a discrete-time residual transformation
@@
The common optimization paradigm in learning neural function approximations require a sufficiently large dataset to prevent overfitting, especially for highly nonlinear dynamics. When data is scarce, alternative strategies are necessary. Classical system identification methods  (Voss et al., 2004) estimate structured model parameters using statistical techniques. Meanwhile, hybrid neural system identification approaches (Chen et al., 2021; Forgione and Piga, 2021) incorporate inductive biases, such as physics-based constraints or regularized latent representations, to improve generalization. To obtain the gradient of the objective w.r.t. the parameters ${\bm{\theta}}$, it is required to propagate the gradients through the numerical integration. A straightforward approach is to store all of the function evaluations in the forward pass of the ODE and backpropagate through the whole numerical integration. Yet, the space complexity increases linearly with time $T$, which becomes a computational burden for long-time series and high-order numerical solvers (Norcliffe and Deisenroth, 2023). Given a scalar-valued loss function $L$, the adjoint of ${\bm{x}}(t)$, ${\bm{a}}(t)={\textnormal{d}}L/{\textnormal{d}}{\bm{x}}(t)$, having the dynamic derived from the chain rule,
@@
As the gradient is obtained as the solution of the time-reversed ODE, the gradients can be evaluated without storing any additional data. Thus, the adjoint method optimizes computational resources by balancing time against space complexity. Several works (Gholaminejad et al., 2019; Zhuang et al., 2020; Onken and Ruthotto, 2020) showed that the gradient estimation based on the adjoint method is error-prone due to the numerical inaccuracies of solving the backward ODE, due to the typical high Lyapunov characteristic exponent of highly nonlinear ODEs. Many improved backward integrators have been proposed to mitigate this issue (Zhuang et al., 2020; 2021a; Gholaminejad et al., 2019). Concurrently, Ott et al. (2020) studied the validity of neural ODEs representing continuous flows, as true continuous-time analogs of ResNets. They claimed that valid neural ODE is invariant to solver configurations, achievable through critical step-size adaptations for integrators. Krishnapriyan et al. (2023) then generalized neural ODE architecture to RK4 integration scheme, showing generalization and convergence to true continuous flow.
@@
Augmented neural ODEs allow observed trajectories to intersect and enable the learning of more general dynamics over the observed state space. Furthermore, according to Dupont et al. (2019), augmented neural ODEs can achieve lower losses, better generalization, and lower computational cost than regular neural ODEs, also since time discretizations can be coarser as the learned rate functions ${\bm{f}}_{\bm{\theta}}$ are smoother. However, challenges are the determination of the size of ${\bm{h}}$ and training without knowing the values of ${\bm{h}}$.
@@
We highlight that the adjoint method has been used in the context of gradient-based shape optimization (Zahr and Persson, 2016). However, utilizing the adjoint method for learning PDE is unexplored. Earliest works propose to learn a convolutional neural network as a map between finite Euclidean spaces $f_{{\bm{\theta}}}$ (Guo et al., 2016; Zhu and Zabaras, 2018; Bhatnagar et al., 2019). Building on these approaches, auto-regressive methods have been introduced to estimate the solution at discrete time steps ${{\bm{u}}({\bm{x}},t+\Delta t)=f_{\bm{\theta}}(\Delta t,{\bm{u}}({\bm{x}},t))}$, and either, mix with conventional numerical methods in classical PDE solvers (Bar-Sinai et al., 2019; Greenfeld et al., 2019; Hsieh et al., 2019), or propagate the solution through time by recurrent calls of the networks (Brandstetter et al., 2022). All previous learning PDE methods have in common that they rely on a discretization of the domain, yet in many scenarios, it is desirable to evaluate the solution at any point in the domain. For these reasons, learning approaches that directly learn a mapping from a point in the domain to its solution $u_{{\bm{\theta}}}$ have been actively explored in recent years (Yu et al., 2018; Raissi et al., 2019; Bar and Sochen, 2019), which introduces physics constraints as losses, is covered in detail in section 3.3. Yet, these models come at the cost that they provide solutions for a particular PDE instance and do not generalize to different inputs. Addressing these contemporary issues of learning PDEs, directly learning the solution operator of the PDE as a map between infinite-dimensional spaces has gained traction in recent years (Lu et al., 2019; Li et al., 2020a; 2021a; Gupta et al., 2021). These models extend the fundamental work of Chen and Chen (1995) that first showed a universal approximation theorem for learning operator networks. These models are explored in more detail in section 3.4.
@@ 3.2 Learning Models with Algebraic Structures
where ${\bm{\tau}}$ is the generalized force. Liu et al. (2005) optimize directly the system trajectory and dynamic parameters, minimizing the violation of the constraint from equation (13). Much later on, by realizing the rigid-body kinetic energy ${\mathcal{T}}({\bm{q}},\dot{{\bm{q}}})=\frac{1}{2}\dot{{\bm{q}}}^{\intercal}{\bm{M}}({\bm{q}})\,\dot{{\bm{q}}}$ under gravitational potential field ${\textnormal{d}}{\mathcal{V}}/{\textnormal{d}}{\bm{q}}={\bm{g}}({\bm{q}})$, Lutter and Peters (2023) propose deep Lagrangian networks (DeLAN), parameterizing directly the inertia tensor ${\bm{M}}_{{\bm{\theta}}}({\bm{q}})$ with positivity constraint ensuring dynamical stability property, and the gravity force ${\bm{g}}_{{\bm{\theta}}}({\bm{q}})$ (Figure 2). Combining automatic differentiation with the state partial derivatives, the derived derivatives w.r.t. time and, thus, the violation of Lagrangian constraint are efficiently computed in a forward pass (see Figure 2), effectively facilitating robot model learning with sample efficiency. Cranmer et al. (2020b) then generalizes DeLAN to the Lagrangian Neural Network (LNN), which assumes a more general form of kinetic energy rather than rigid-body kinetics. The LNN works by directly computing the matrix inverse of the Hessian of a learned Lagrangian $\left(\frac{\partial^{2}{\mathcal{L}}}{\partial\dot{{\bm{q}}}^{\top}\partial\dot{{\bm{q}}}}\right)$, and computing the implied acceleration
@@
where the inverse is typically approximated with a Moore-Penrose pseudoinverse. However, employing the Moore-Penrose pseudoinverse may cause numerical instability due to ill-conditioned Hessians and suffers from poor scalability in high dimensions (Golub and Van Loan, 2013). The training is then done via backpropagating through a loss between the true and predicted $\ddot{{\bm{q}}}$.
@@
with neural networks on the damping ${\bm{D}}$ matrix, the Hamiltonian ${\mathcal{H}}$, the generalized force ${\bm{G}}({\bm{q}}){\bm{u}}$, effectively recovering the underlying stationary Hamiltonian, time-dependent force, and dissipative coefficient. Recently, Roth et al. (2025) introduced the concepts of Lyapunov and global stability of dynamic systems into Port-Hamiltonian NNs by enforcing the Hamiltonian to be convex and possess a strict minimum. This additional physical bias and mathematical structure greatly improve robustness of training and predictions as instabilities are avoided, enhance extrapolation capabilities, and reduce model variance.
@@
where ${\bm{q}}(t)\in{\mathbb{R}}^{d}$ are the dynamic states and ${\bm{z}}(t)\in{\mathbb{R}}^{n}$ are the algebraic variables. This class of differential-algebraic equations has been used to model various engineering applications, such as power network systems with frequency-dependent dynamic loads or nonlinear oscillations. Under the AINN framework, Moya and Lin (2023) model $f,g$ as neural networks and treat both statements in equation (18) as algebraic relations. Given training data on ${\bm{q}}$, they perform a model rollout with an implicit Runge-Kutta method (Iserles, 2009) from the initial conditions, then learn $f,g$ by matching the rollout trajectories against the training data and minimizing the violation of the constraint $g({\bm{q}},{\bm{z}})$, satisfying the defined algebraic relations.
AINN is still in its infancy state with very few works. However, we foresee vast applications of parameterizing algebraic rules with neural networks in many hybrid systems (i.e., discrete-continuous) in robotics or engineering applications.
3.3 Learning Simulation Solutions with Physics-Informed Losses
where ${\bm{v}}_{m}$ are test functions and $M$ is the number of sampling points. This approach is attractive because it can lower the order of differential operators that the neural network approximates, making the computation of loss functions cheaper. However, the analytic tractability puts severe restrictions on the models and settings that can be considered, and the authors only demonstrate the approach on a one- and two-dimensional case of Poisson’s equation.
The practical challenges of applying PINNs in dynamical systems lie in approximating and optimizing the loss function. Due to the nature of differential equations, the residual loss is often not ’fixed’ between time steps, as it requires the model to be self-consistent with a set of its own derivatives (Wang et al., 2021a) . This objective means the regression target is not fixed, and the model is regressed against itself during optimization until self-consistency is achieved. This ‘bootstrapping’ of a function approximator can lead to unstable optimization in practice [^2] Bootstrapping and function approximation comprises of two parts of the ‘deadly triad’ in reinforcement learning (Sutton and Barto, 2018) where this model bootstrapping is also performed when computing value functions. As a result, sampling and optimization must be designed carefully and jointly to ensure successful optimization.
Subramanian et al. (2023) propose an adaptive self-supervision algorithm to enhance the performance and ‘trainability’ of PINNs. By reallocating collocation points to regions with higher errors and periodically resampling them, the proposed methods effectively reduce prediction errors and address optimization challenges across various problems, without increasing computational costs.
Optimization. Wang et al. (2021b) investigate the training dynamics of PINNs and observed that the magnitude of the parameter gradient increased during training, in part due to the multi-objective loss function. To achieve stable convergence, the authors advocated for a learning rate schedule.
Fonseca et al. (2023) studied the optimization landscape of PINNs, and found that their solutions lie in sharp minima, i.e., regions of the loss landscape where the loss varies significantly under small perturbations of the solution. This is in contrast to folklore in deep learning, where solutions in flat minima are preferred as they are interpreted as solutions that are not overfitting and, therefore, generalizing. Since PINNs models are trained in their domain, the issue of out-of-distribution generalization is less of a concern. This difference results in typical deep learning optimization techniques being less desirable. The authors found that quasi-Newton solvers such as the limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm (LBFGS) were superior optimizers due to their strength in reaching sharp minima by incorporate loss curvature.
@@ 3.4 Operator Learning
So far, this survey discussed learning the solution $u_{\bm{\theta}}({\bm{x}})$ to a differential equation for a specific configuration ${\bm{a}}$, (possible) boundary conditions (equation (4)), and source terms ${\bm{f}}$. However, this approach requires solving the system anew for each new configuration. A more efficient alternative is to learn an operator — a model that maps an infinite-dimensional function-space input to the corresponding solution. This enables generalization across different configurations without solving the differential equation from scratch. Specifically, we introduce learning a solution operator ${\mathsf{L}}_{\bm{a}}^{-1}{\bm{f}}({\bm{x}})={\bm{u}}_{\bm{a}}({\bm{x}})$ that maps from state-varying functions ${\bm{a}}({\bm{x}}):\Omega\rightarrow{\mathbb{R}}^{d_{a}}$ to the corresponding solution (Figure 4). Operating on functions introduces the same flexible discretization invariance discussed in section 3.1, but now extended beyond time to any continuous state space. Moreover, the training data for such models now extends across several instances of the physical system, so the model is extended in a physics-informed fashion to be conditioned on the system parameters. We will revisit multi-instance learning in section 4 in the context of learning data-driven priors.
@@
Kovachki et al. (2023) report that training with the relative $L^{2}$ loss equation (23) is less prone to overfitting due to normalization compared to the alternative mean squared error (MSE) loss. The specific parameterization of the operator networks has been the primary focus of recent approaches to guarantee discretization invariance, which has a bounded generalization error while being computationally efficient. Following this, we will discuss the two prevailing operator learning architectures, DeepONets (Lu et al., 2019) and neural operators (Li et al., 2020a) highlighted in Figure 4. Although both architectures follow universal approximation theorems to learn arbitrary operators, their conceptual design is fundamentally different.
Figure 4: (Left) A schematic of operator networks that learn a functional mapping from the inputs $a$ to the PDE solution ${\bm{u}}$ at a designated point ${\bm{x}}$. (Middle) DeepONets can approximate any nonlinear operator by learning a finite-dimensional mapping $\hat{\bm{u}}={\bm{\varphi}}_{\bm{\theta}}(\hat{\bm{a}})$ in a latent space which is spanned by linear encodings $\hat{\bm{a}}={\bm{e}}_{\bm{\theta}}({\bm{a}})$ and linear decoding ${\bm{u}}={\bm{g}}_{\bm{\theta}}(\hat{\bm{u}},{\bm{x}})$. (Right) Neural operators directly map a discretized representation of the input ${\bm{a}}_{1:N}$ to the solution at an arbitrary discretization ${\bm{u}}_{1:M}$.
@@
In order to deal with infinite-dimensional input and output spaces, deep operator networks (DeepONets) (Lu et al., 2019) are based on the idea of learning the solution operator in a finite-dimensional space ${\bm{\varphi}}:{\mathbb{R}}^{d_{{\bm{e}}}}\rightarrow{\mathbb{R}}^{d_{{\bm{g}}}}$. This requires a functional mapping of the input functions ${\bm{a}}$ into a latent representation ${\bm{e}}:{\mathcal{A}}\rightarrow{\mathbb{R}}^{d_{{\bm{e}}}}$ and a respective functional mapping ${\bm{g}}:{\mathbb{R}}^{d_{{\bm{g}}}}\rightarrow{\mathcal{U}}$ that decodes. Thus, the operator learning architecture based on the finite-dimensional mappings can be formulated as ${\mathsf{S}}_{\bm{\theta}}{\bm{a}}={\bm{g}}_{\bm{\theta}}\circ{\bm{\varphi}}_{\bm{\theta}}\circ{\bm{e}}_{\bm{\theta}}$. For this architecture, Chen and Chen (1995) first showed that for linear maps ${\bm{e}}_{\bm{\theta}}$ and ${\bm{g}}_{\bm{\theta}}$, and a single-layer network ${\bm{\varphi}}_{\bm{\theta}}$, there exist finite-dimensional latent spaces ${\mathbb{R}}^{d_{\bm{e}}}$ and ${\mathbb{R}}^{d_{\bm{g}}}$ for which any nonlinear operator can be approximated. This universal approximation theorem for operator learning has recently been expanded to encompass arbitrary deep-learning architectures for mapping in finite-dimensional spaces (Lu et al., 2019). The different embodiments of the three modules lead to distinct approaches, all underlying the same concept of finite-dimensional mappings.
@@
Similar ideas within the framework of reduced-order methods have been presented by Peherstorfer and Willcox (2016). This method, known as operator inference, involves projecting the state-space onto a lower-dimensional representation using proper orthogonal decomposition (POD) and subsequently learning the linear operators of systems of nonlinear ordinary differential equations (ODEs). However, these models often lack the generalization capabilities seen in more recent methodologies, such as neural operators.
@@
In numerous relevant applications, the PDE operator exhibits nonlinearity, rendering equation (24) inapplicable. Although approaches utilizing encoding and decoding schemes to transform the problem into sub-spaces amenable to linear operators have been suggested (Gin et al., 2021), they lack the theoretical assurances inherent in operator learning methods. Yet, the idea of using integral kernel operators play a pivotal role in designing neural operators with theoretical guarantees
@@
The kernel integration layers are composed of a linear operator, represented by the matrix ${\bm{W}}_{t}\in{\mathbb{R}}^{d_{v_{t}}\times d_{v_{t-1}}}$, a bias term ${\bm{b}}_{t}\in{\mathbb{R}}^{d_{v_{t}}}$, and the integral kernel operator (equation (25)). By parameterizing ${\mathsf{P}},{\mathsf{Q}},{\bm{W}}_{t},{\bm{b}}_{t},\kappa_{t}$, the neural operator can be shown to approximate arbitrary nonlinear operators (Kovachki et al., 2023). Although NOs are not restricted to finite-dimensional mappings, as noted by (Kovachki et al., 2024), in practice, NOs work on discretized representations of the input domain ${\bm{a}}_{1:N}$. The main reason is that the integral kernel operator (equation (25)) remains a computationally demanding operation. Therefore, we will examine different architectures to approximate ${\mathsf{K}}$ efficiently.